# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v3

@@ III.2 A common language to discuss
The underlying goal of a quantum benchmark is to be able to compare devices with each other. What is more, similarly to classical benchmarking, quantum benchmarks involve running specific tests or workloads on the systems and software collecting data to analyze its performance. Examples of quantum benchmarks include QASMBench li2022qasmbench , Quantum Volume PhysRevA.100.032328 , or Quantum LINPACK QuantumLINPACK .
@@
When the contextual information provides the details about the system being assessed, we briefly refer to them as just metrics. Interestingly, Quantum Volume is not only a benchmark but also a metric PhysRevA.100.032328 , even though it is tightly coupled to the Quantum Volume benchmark definition. Other examples of metrics are cross-entropy difference, $\ell_{1}$-Norm Distance to measure total divergence of probability distributions to the sample space, the gates fidelity, the number of qubits, or the Q-Score Martiel2021BenchmarkingQC .
@@
This concept is particularly relevant, but it comes with challenges. In particular, if the approach to verify correctness lays on the assumption that the device would be able to execute correctly the output of certain input instances, we need to be able to answer the following question Gheorghiu2018 : ”If a quantum experiment solves a problem which is proven to be intractable for classical computers, how can one verify the outcome of the experiment?”. Or, as the Aaronson’s \$25.00 prize states AaronsonPrize : ”If a quantum computer can efficiently solve a problem, can it also efficiently convince an observer that the solution is correct? More formally, does every language in the class of quantumly tractable problems (BQP) admit an interactive proof where the prover is in BQP and the verifier is in the class of classically tractable problems (BPP)?”. Unfortunately, no one has claimed the prize yet. Still, there are other approaches worth flagging in the field leveraging quantum provers as trusted measurement devices Mahadev2018 .
@@
In this category, we can find approaches such as the circuit mirroring Proctor2021 , or quantum state tomography PhysRevA.40.2847 , which can be leveraged to help quantum benchmark designers when trying to measure the results of their computations.
We need to clarify that quantum verification and the quantum benchmarking verification techniques are different from the classic concept of formal verification of programs. Indeed, there exists a parallel research field dedicated to the formal verification of quantum programs FormalVerificationQuantumPrograms , and we can easily extend the notion from their classic counterparts. Note that even though the formal verification of programs is an interesting field with intensive research for both classical and quantum cases, historically a testing-based approach that does not rely on any mathematical formalization for the proof of correctness has been favoured. For those cases, we introduce the idea of testing quantum programs:
@@
There are some early works in the field of testing GarcadelaBarrera2021 . Some of them raise awareness about the challenges of introducing classical computing techniques for testing such as white- and black-box testing 8805685 , and hence the need for an enhanced approach for quantum computers. The importance of quantum testing will become greater in the upcoming years as we transition to the next eras.
@@
The most well-known example in this category would be the volumetric benchmark framework Blume_Kohout_2020 .
@@
III.3 Reviewing existing proposals for benchmarking
@@
Recently, there have been attempts to categorize benchmarking practices. However, the lack of a precise definition of concepts has led to confusion in the field of quantum benchmarking. There are instances where the concepts of certification and benchmarking are intertwined within a shared framework eisert2020quantum , or concepts such as metrics and benchmarks are sometimes used interchangeably SoKBenchmarkingQuantumComputer . In Ref. SoKBenchmarkingQuantumComputer , the term physical benchmarkings has been introduced to describe properties of a quantum processor that are more accurately understood as metrics. Additionally, the concept of aggregated benchmarks has been presented, which combines metrics and benchmarking techniques. For instance, Circuit Layer Operations per Second https://doi.org/10.48550/arxiv.2110.14108 is a metric while Mirroring Circuits Proctor2021 ) is a technique employed in benchmarks. Other proposal more aligned with the intent of distinguishing between different terms have been presented proctor2024benchmarkingquantumcomputers . An excellent example of such studies is the article by Resch and Karpuzcu on the impact of quantum noise https://doi.org/10.48550/arxiv.1912.00546 .
@@
On the other hand, in Ref. https://doi.org/10.48550/arxiv.2110.14108 the researchers from the IBM T. J. Watson Research Center took a different approach. They introduced the benchmarking Pyramid which classified in three levels the different benchmarks: Holistic (measuring the systems as a whole), Subsystem (measuring different aspects of the subcomponents used to assemble the system, e.g. classical control logic system), and Device (e.g. quantum gates or qubit registers). Apart from classifying them into these three level groups, they also distinguished two main factors that can be measured for each level: Quality (a property that determines the size of quantum circuits that can be faithfully executed), and Speed (a property related to the number of circuits that the quantum computer system can execute per unit of time). The hierarchization process carried out is interesting, and it accounts for a division of different types of, what we call, metrics. Still, during their work they focused on a procedure to generate a particular test for Quantum Volume, later defining metrics such as CLOPS on top of it but limiting the scope of the benchmark. Indeed, there are well-known extensions to the Quantum Volume to build a family of volumetric benchmarks that generalise these types of tests Blume_Kohout_2020 . Other remarkable attempts to build thorough benchmarks can be found in the literature, such as the Application-Oriented Performance Benchmarks from the Quantum Economic Development Consortium (QED-C) quantum_consortium ; lubinski2023applicationoriented , the proposal from Daniel et al. Mills2021 where they proposed measuring the performance based on three types of circuits (deep, shallow, and square classes) with three metrics that can be computed with classical computers (heavy output generation probability, cross-entropy difference, and $\ell_{1}$-Norm Distance), or the QUARK framework Finzgar_2022 .
@@ III.3.1 Techniques
Mirror circuits: The concept of mirror circuits Proctor2021 ; PhysRevLett.129.150502 , inspired by Loschmidt echo, involves creating an alternative efficient classically simulatable circuit that mirrors a given quantum circuit $C$. The mirror circuit technique generally consists of implementing a circuit $C$ followed by its inverse $C^{-1}$. There are several variations or specifications of this method. One common approach involves preparing the qubits in randomized Pauli eigenstates, executing the base circuit $C$, applying a randomly selected Pauli layer $Q$, running the reversed circuit $C^{-1}$ adjusted to implement a specific Pauli operation, and finally measuring the qubits in their initial basis. Another variation lies in how $C^{-1}$ is implemented, either by reversing all the gates in $C$, or by applying a different random transpilation of the circuit.
@@
Classical shadow tomography: Classical shadow tomography addresses challenges in efficient information extraction from quantum states doi:10.1137/18M120275X ; article12 ; zhou2023efficient . Traditional methods demand an exponential number of measurements for a complete description of $N$-qubit states, which is impractical as discussed in Section III.1. However, classical shadow tomography provides an efficient alternative for extracting information with a limited number of measurements when detailed knowledge of the entire density matrix is unnecessary. This technique uses randomized measurements zhou2023efficient , employing independent random unitary operations, $U$, selected from an ensemble. Before measurement, this operation is applied, yielding a classical snapshot $\sigma_{U,b}=U^{\dagger}|b\rangle\langle b|U$ with a probability $p_{U,b}$. The average snapshot $\sigma$ relates to the original density matrix $\rho$ through a quantum channel $M$ as $\sigma=\sum_{b}\mathbb{E}_{\mathcal{E}_{u}}[\sigma_{U,b}p_{U,b}]=M(\rho)$. Physical observables, denoted as $\langle O\rangle$, can be expressed as $\langle O\rangle=\text{tr}\left(OM^{-1}(\sigma)\right)$. Consequently, efficient classical computation of $M^{-1}$ and a low sampling complexity are key criteria for effective classical shadow tomography schemes. Different schemes correspond to diverse choices of the ensemble $\mathcal{E}_{U}$. For instance, utilizing Haar random unitaries and tensor products of single-qubit Haar random unitaries, one can prove that while the shadow norm $\|O\|^{2}_{\text{sh}}$ scales as $2^{N}$ for general $N$-qubit systems, it scales as $3^{k}$ for Pauli operators with a size of $k$, showing a better performance for few-body operators.
@@
Efficient classically simulatable quantum circuits: Several families of quantum circuits can be efficiently simulated in classical computers. Particularly relevant are Clifford circuits, a specific class of quantum circuits composed entirely of Clifford gates. These gates include the Hadamard gate ($H$), the Phase gate ($S=\sqrt{Z}$), and the Controlled-NOT gate (CNOT). The efficient simulatability of these circuits is ensured by Gottesman-Knill theorem, which states that any quantum circuit comprising an initial state in the computational basis, Clifford gates, and measurement in the computational basis, can be efficiently simulated in a classical computer 10.5555/863284 ; nielsen00 .
@@
Randomized Benchmarking (RB): Randomized Benchmarking (RB) is a technique used for characterizing the error rate of quantum processors. This technique involves applying a sequence of random gates, which can be sampled uniformly at random from the Haar distribution Emerson_2005 or uniformly at random from any unitary 2-design Dankert_2009 , such as the Clifford group. RB has been employed as a technique for measuring errors, such as the gate fidelity of Clifford gates PhysRevLett.106.180504 , the coherent or incoherent nature of the noise affecting Clifford gates Wallman_2015 , or the noise of a particular gate Magesan_2012 , among others. This technique is robust against state preparation and measurement (SPAM) noise and is considered scalable. Although the standard randomized benchmarking (RB) protocol employing this technique is theoretically scalable, in practice it becomes challenging to implement for more than 5 or 6 qubits due to the depth of the required Clifford circuits. To overcome this limitation, recent variants of the protocol have been developed that aim to minimize circuit depth, making the method viable for systems with up to 20 qubits Hines2024 ; PhysRevLett.129.150502 ; polloreno2023 . For a detailed review of the different types of protocols based on this technique, see hashim2024 .
III.3.2 Metrics
@@
Let us introduce some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer. We will categorize these metrics into two types: Static and Non-static metrics.
Static metrics refer to internal features of the quantum processor that define specific performance-related properties. Once the processor is manufactured, these metrics cannot be altered. Next, we introduce the main static metrics:
@@
Number of qubits: One of the most commonly used metrics in quantum performance evaluation is the number of (working connected) qubits, which directly impacts the quantum information processing capacity https://doi.org/10.48550/arxiv.2110.14108 . As the number of qubits increases, the potential of the quantum computer to address larger and more complex problems grows exponentially. We emphasize the importance of considering not only the quantity but also the quality of qubits, which is especially relevant in the NISQ era. Indeed, scaling quantum computers to a large number of qubits is a challenging engineering task that involves addressing various technical issues depending on the characteristics of the physical platform in use.
@@
Qubits connectivity: Qubit connectivity is an important metric to consider because it determines the direct interactions between the qubits Holmes_2020 . Quantum algorithms require the interaction of multiple qubits through quantum gates. When qubits are not directly connected additional quantum gates could be necessary to facilitate their interaction increasing the runtime of the algorithm, circuit depth, and the risk of errors in the NISQ era. Therefore, qubit connectivity directly affects the performance of the quantum computer. Currently, physical architectures provide varying qubit connectivity options, such as nearest-neighbor connectivity Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 or fully connected connectivity QCion ; Zhang_2017 . One could provide the maximal, minimal and average connectivity, for instance. Other method of mapping the qubit connectivity into a numerical representation involves constructing a $n\times n$ matrix of normalized coupling terms for a $n$-qubit processor and compute its norm. This approach has as an advantage that it takes into account also the quality of the couplings but the norm complicates comparing processors with different sizes.
@@
Gate fidelity: The gate fidelity refers to the accuracy of quantum operations. There is no unique definition of fidelity for a quantum operation; however, most definitions are linearly related and share similar properties. Therefore, throughout this text, we refer to fidelity in a general sense. In practical applications, however, it is important that users specify the particular fidelity measure being used. The gate fidelity focuses on how well a quantum gate performs the intended operation compared to the ideal outcome. When a quantum gate is applied, there is a predetermined ideal output, but due to various factors like noise and imperfections in the quantum processor, the actual observed output might deviate from the expected one. High gate fidelity is essential for the reliability and precision of computations. If the gate fidelity is low, it indicates that the quantum gate is introducing errors, reducing the accuracy of quantum algorithms. However, although the concept is clear, its practical application in a quantum processor presents some complex problems. In quantum platforms where entangling gates are natively implemented by two-qubit gates, such as in superconducting circuits or photonic quantum computing, it is common practice to characterize the performance of quantum hardware using metrics associated with two-qubit gates. In such cases, it is natural to evaluate the average, maximum, or minimum fidelity of these gates. However, this logic does not directly apply to other platforms, such as trapped-ion or cold-atom systems, where native multi-qubit gates can be implemented. In these architectures, a single operation can generate collective entanglement among three or more qubits without the need to decompose the operation into a sequence of two-qubit gates. This capability fundamentally change the structure of quantum computation and the efficiency with which certain algorithms can be executed. While comparing individual two-qubit gate fidelities across platforms can be valid when the same operations are implemented, relying solely on such metrics to assess the overall performance of a quantum processor can be misleading. This is particularly true when native gate sets differ significantly, as is the case between platforms supporting multi-qubit gates and those limited to two-qubit gates. In this case, moreover, the fidelity is substantially more costly to calculate. It is also important to note that gate fidelity is not a static property of a quantum hardware. It can vary over time due to coherence times of qubits that are highly sensitive to environmental. Additionally, the fidelity depends on implementation details such as pulse shaping and calibration routines, which can be optimized without manufacture a new quantum hardware.
Readout fidelity: The readout fidelity is a metric that quantifies the accuracy with which a quantum processor measures the final state of qubits. It reflects the probability that the measurement outcome corresponds to the quantum state, typically in the computational basis. High readout fidelity is essential for reliable data collection and accurate measurement of qubit states, especially in NISQ devices, where measurement errors can significantly impact overall performance. Importantly, several techniques can be employed to improve readout fidelity without the need to manufacture a new quantum processor PRXQuantum.5.040326 ; PhysRevApplied.23.054057 ; Seif_2018 .
Decoherence times: The decoherence times refer to the period during which the qubit can maintain its quantum coherence. The interaction of qubits with the environment results in the loss of quantum superposition and the decay from the excited state to the ground state, characterized by the times $T_{2}$ and $T_{1}$ respectively. Longer decoherence times are preferable as they might permit deeper circuits to be run, reducing the probability of errors and improving computation accuracy. Unfortunately, coherence times by themselves are not an accurate measure of the quality of the processor. Decoherence occurs due to the coupling of the quantum system to the environment. Strongly protecting a quantum system can result in reduced manipulability and a substantial increase in the time required to operate quantum gates. Consequently, if we double the decoherence times but also double the gate operation times, this has no practical consequence on the processor’s performance, as the number of operations we can carry out will not change. It is important to note that coherence times are not fixed characteristics of quantum hardware, but time-dependent parameters that may vary due to environmental influences. Additionally, techniques such as dynamical decoupling Viola1998 can be used to improve coherence times by mitigating specific sources of noise applying control pulses to qubits.
@@
Gate speeds: Gate speed is a metric that measures how rapidly a quantum computer can execute a given quantum gate van_der_Schoot_2023 . Considering decoherence times, this metric allows us to determine the number of gates that can be performed before the qubits lose coherence. By increasing gate speeds, a greater number of gates can be executed within the same time interval, leading to more applied gates during the lifetime of the qubits. Here, the argument explained for coherence times can be inverted, since fast manipulable systems could also be susceptible to decoherence. The problems of this metric are similar or worse to gate fidelity metric. It is important to highlight that gate speed is not a fixed metric, as it can be optimized through pulse engineering techniques Werninghaus2021 . Moreover, quantum computing platforms, the strength of the interaction between qubits can be dynamically tuned via external control parameters Yan2018 . Therefore, both pulse engineering techniques and dynamic tuning of qubit interactions provide control over quantum gate speeds.

### v4

@@ 3.3 A common language to discuss
The underlying goal of a quantum benchmark is to be able to compare devices with each other. What is more, similarly to classical benchmarking, quantum benchmarks involve running specific tests or workloads on the systems and software collecting data to analyze its performance. Examples of quantum benchmarks include QASMBench [96], Quantum Volume [97], or Quantum LINPACK [98].
@@
When the contextual information provides the details about the system being assessed, we briefly refer to them as just metrics. Interestingly, Quantum Volume is not only a benchmark but also a metric [97], even though it is tightly coupled to the Quantum Volume benchmark definition. Other examples of metrics are cross-entropy difference, $\ell_{1}$-Norm Distance to measure total divergence of probability distributions to the sample space, the gates fidelity, the number of qubits, or the Q-Score [99].
@@
This concept is particularly relevant, but it comes with challenges. In particular, if the approach to verify correctness lays on the assumption that the device would be able to execute correctly the output of certain input instances, we need to be able to answer the following question [100]: “If a quantum experiment solves a problem which is proven to be intractable for classical computers, how can one verify the outcome of the experiment?”. Or, as the Aaronson’s \$25.00 prize states [101]: “If a quantum computer can efficiently solve a problem, can it also efficiently convince an observer that the solution is correct? More formally, does every language in the class of quantumly tractable problems (BQP) admit an interactive proof where the prover is in BQP and the verifier is in the class of classically tractable problems (BPP)?”. Unfortunately, no one has claimed the prize yet. Still, there are other approaches worth flagging in the field leveraging quantum provers as trusted measurement devices [102].
@@
In this category, we can find approaches such as the circuit mirroring [103], or quantum state tomography [104], which can be leveraged to help quantum benchmark designers when trying to measure the results of their computations.
We need to clarify that quantum verification and the quantum benchmarking verification techniques are different from the classic concept of formal verification of programs. Indeed, there exists a parallel research field dedicated to the formal verification of quantum programs [105], and we can easily extend the notion from their classic counterparts. Note that even though the formal verification of programs is an interesting field with intensive research for both classical and quantum cases, historically a testing-based approach that does not rely on any mathematical formalization for the proof of correctness has been favoured. For those cases, we introduce the idea of testing quantum programs:
@@
There are some early works in the field of testing [106]. Some of them raise awareness about the challenges of introducing classical computing techniques for testing such as white- and black-box testing [107], and hence the need for an enhanced approach for quantum computers. The importance of quantum testing will become greater in the upcoming years as we transition to the next eras.
@@
The most well-known example in this category would be the volumetric benchmark framework [108].
@@
3.4 Reviewing existing proposals for benchmarking
@@
Recently, there have been attempts to categorize benchmarking practices. However, the lack of a precise definition of concepts has led to confusion in the field of quantum benchmarking. There are instances where the concepts of certification and benchmarking are intertwined within a shared framework [109], or concepts such as metrics and benchmarks are sometimes used interchangeably [110]. In Ref. [110], the term physical benchmarkings has been introduced to describe properties of a quantum processor that are more accurately understood as metrics. Additionally, the concept of aggregated benchmarks has been presented, which combines metrics and benchmarking techniques. For instance, Circuit Layer Operations per Second [111] is a metric while Mirroring Circuits [103]) is a technique employed in benchmarks. Other proposal more aligned with the intent of distinguishing between different terms have been presented [95]. An excellent example of such studies is the article by Resch and Karpuzcu on the impact of quantum noise [112].
@@
On the other hand, in Ref. [111] the researchers from the IBM T. J. Watson Research Center took a different approach. They introduced the benchmarking Pyramid which classified in three levels the different benchmarks: Holistic (measuring the systems as a whole), Subsystem (measuring different aspects of the subcomponents used to assemble the system, e.g. classical control logic system), and Device (e.g. quantum gates or qubit registers). Apart from classifying them into these three level groups, they also distinguished two main factors that can be measured for each level: Quality (a property that determines the size of quantum circuits that can be faithfully executed), and Speed (a property related to the number of circuits that the quantum computer system can execute per unit of time). The hierarchization process carried out is interesting, and it accounts for a division of different types of, what we call, metrics. Still, during their work they focused on a procedure to generate a particular test for Quantum Volume, later defining metrics such as CLOPS on top of it but limiting the scope of the benchmark. Indeed, there are well-known extensions to the Quantum Volume to build a family of volumetric benchmarks that generalise these types of tests [108]. Other remarkable attempts to build thorough benchmarks can be found in the literature, such as the Application-Oriented Performance Benchmarks from the Quantum Economic Development Consortium (QED-C) [113, 114], the proposal from Daniel et al. [115] where they proposed measuring the performance based on three types of circuits (deep, shallow, and square classes) with three metrics that can be computed with classical computers (heavy output generation probability, cross-entropy difference, and $\ell_{1}$-Norm Distance), or the QUARK framework [116].
@@ 3.4.1 Techniques
Mirror circuits: The concept of mirror circuits [103, 117], inspired by Loschmidt echo, involves creating an alternative efficient classically simulatable circuit that mirrors a given quantum circuit $C$. The mirror circuit technique generally consists of implementing a circuit $C$ followed by its inverse $C^{-1}$. There are several variations or specifications of this method. One common approach involves preparing the qubits in randomized Pauli eigenstates, executing the base circuit $C$, applying a randomly selected Pauli layer $Q$, running the reversed circuit $C^{-1}$ adjusted to implement a specific Pauli operation, and finally measuring the qubits in their initial basis. Another variation lies in how $C^{-1}$ is implemented, either by reversing all the gates in $C$, or by applying a different random transpilation of the circuit.
@@
Classical shadow tomography: Classical shadow tomography addresses challenges in efficient information extraction from quantum states [118, 119, 120]. Traditional methods demand an exponential number of measurements for a complete description of $N$-qubit states, which is impractical as discussed in Section 3.1. However, classical shadow tomography provides an efficient alternative for extracting information with a limited number of measurements when detailed knowledge of the entire density matrix is unnecessary. This technique uses randomized measurements [120], employing independent random unitary operations, $U$, selected from an ensemble. Before measurement, this operation is applied, yielding a classical snapshot $\sigma_{U,b}=U^{\dagger}|b\rangle\langle b|U$ with a probability $p_{U,b}$. The average snapshot $\sigma$ relates to the original density matrix $\rho$ through a quantum channel $M$ as $\sigma=\sum_{b}\mathbb{E}_{\mathcal{E}_{u}}[\sigma_{U,b}p_{U,b}]=M(\rho)$. Physical observables, denoted as $\langle O\rangle$, can be expressed as $\langle O\rangle=\text{tr}\left(OM^{-1}(\sigma)\right)$. Consequently, efficient classical computation of $M^{-1}$ and a low sampling complexity are key criteria for effective classical shadow tomography schemes. Different schemes correspond to diverse choices of the ensemble $\mathcal{E}_{U}$. For instance, utilizing Haar random unitaries and tensor products of single-qubit Haar random unitaries, one can prove that while the shadow norm $\|O\|^{2}_{\text{sh}}$ scales as $2^{N}$ for general $N$-qubit systems, it scales as $3^{k}$ for Pauli operators with a size of $k$, showing a better performance for few-body operators.
@@
Efficient classically simulatable quantum circuits: Several families of quantum circuits can be efficiently simulated in classical computers. Particularly relevant are Clifford circuits, a specific class of quantum circuits composed entirely of Clifford gates. These gates include the Hadamard gate ($H$), the Phase gate ($S=\sqrt{Z}$), and the Controlled-NOT gate (CNOT). The efficient simulatability of these circuits is ensured by Gottesman-Knill theorem, which states that any quantum circuit comprising an initial state in the computational basis, Clifford gates, and measurement in the computational basis, can be efficiently simulated in a classical computer [121, 122].
@@
Randomized Benchmarking (RB): Randomized Benchmarking (RB) is a technique used for characterizing the error rate of quantum processors. This technique involves applying a sequence of random gates, which can be sampled uniformly at random from the Haar distribution [123] or uniformly at random from any unitary 2-design [124], such as the Clifford group. RB has been employed as a technique for measuring errors, such as the gate fidelity of Clifford gates [125], the coherent or incoherent nature of the noise affecting Clifford gates [126], or the noise of a particular gate [127], among others. This technique is robust against state preparation and measurement (SPAM) noise and is considered scalable. Although the standard randomized benchmarking (RB) protocol employing this technique is theoretically scalable, in practice it becomes challenging to implement for more than 5 or 6 qubits due to the depth of the required Clifford circuits. To overcome this limitation, recent variants of the protocol have been developed that aim to minimize circuit depth, making the method viable for systems with up to 20 qubits [128, 117, 129]. For a detailed review of the different types of protocols based on this technique, see [93].
3.4.2 Metrics
@@
Let us introduce some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer. We will categorize these metrics into two types: Static and Non-static metrics. Static metrics refer to internal features of the quantum processor that define specific performance-related properties. Once the processor is manufactured, these metrics cannot be altered.
For each metric, we indicate which properties from 2.3 it satisfies. We do not attempt to quantify the degree of satisfaction of each property, as such grading is difficult to define and not essential for the purposes of this work. Accordingly, our classification records only whether a property is satisfied or not. However, when a property is not satisfied, this may occur to varying degrees and for different reasons, including platform heterogeneity, strong dependence on the experimental context, unfavorable scalability, limited ability of the metric to reliably discriminate device performance, or ambiguities in the metric definition itself, among other factors. Next, we introduce the main static metrics:
@@
Number of qubits: One of the most commonly used metrics in quantum performance evaluation is the number of (working connected) qubits, which directly impacts the quantum information processing capacity [111]. As the number of qubits increases, the potential of the quantum computer to address larger and more complex problems grows exponentially. We emphasize the importance of considering not only the quantity but also the quality of qubits, which is especially relevant in the NISQ era. Indeed, scaling quantum computers to a large number of qubits is a challenging engineering task that involves addressing various technical issues depending on the characteristics of the physical platform in use.When the metric is defined as the number of qubits in a quantum computer that form a connected graph, it trivially satisfies all the properties required of a good metric.
@@
Qubits connectivity: Qubit connectivity is an important metric to consider because it determines the direct interactions between the qubits [130]. Quantum algorithms require the interaction of multiple qubits through quantum gates. When qubits are not directly connected additional quantum gates could be necessary to facilitate their interaction increasing the runtime of the algorithm, circuit depth, and the risk of errors in the NISQ era. Therefore, qubit connectivity directly affects the performance of the quantum computer. Currently, physical architectures provide varying qubit connectivity options, such as nearest-neighbor connectivity [60, 61, 62] or fully connected connectivity [58, 59]. One could provide the maximal, minimal and average connectivity, for instance. Other method of mapping the qubit connectivity into a numerical representation involves constructing a $n\times n$ matrix of normalized coupling terms for a $n$-qubit processor and compute its norm. This approach has as an advantage that it takes into account also the quality of the couplings but the norm complicates comparing processors with different sizes. When considering the average connectivity, the metric is clearly practical and repeatable, as it depends only on fixed quantities that do not vary over time. However, it is not reliable, since performance is influenced not only by the number of connections but also by how those connections are arranged and by the quality of the connections. As a result, a quantum computer with fewer connections may allow for shallower circuits than another system with higher average connectivity. For the same reason, the metric is not linear. Finally, it is consistent, as its definition does not depend on the underlying hardware architecture.
@@
Gate fidelity: The gate fidelity refers to the accuracy of quantum operations. There is no unique definition of fidelity for a quantum operation; however, most definitions are linearly related and share similar properties. Therefore, throughout this text, we refer to fidelity in a general sense. In practical applications, however, it is important that users specify the particular fidelity measure being used. The gate fidelity focuses on how well a quantum gate performs the intended operation compared to the ideal outcome. When a quantum gate is applied, there is a predetermined ideal output, but due to various factors like noise and imperfections in the quantum processor, the actual observed output might deviate from the expected one. High gate fidelity is essential for the reliability and precision of computations. If the gate fidelity is low, it indicates that the quantum gate is introducing errors, reducing the accuracy of quantum algorithms. However, although the concept is clear, its practical application in a quantum processor presents some complex problems. In quantum platforms where entangling gates are natively implemented by two-qubit gates, such as in superconducting circuits or photonic quantum computing, it is common practice to characterize the performance of quantum hardware using metrics associated with two-qubit gates. In such cases, it is natural to evaluate the average, maximum, or minimum fidelity of these gates. However, this logic does not directly apply to other platforms, such as trapped-ion or cold-atom systems, where native multi-qubit gates can be implemented. In these architectures, a single operation can generate collective entanglement among three or more qubits without the need to decompose the operation into a sequence of two-qubit gates. This capability fundamentally change the structure of quantum computation and the efficiency with which certain algorithms can be executed. While comparing individual two-qubit gate fidelities across platforms can be valid when the same operations are implemented, relying solely on such metrics to assess the overall performance of a quantum processor can be misleading. This is particularly true when native gate sets differ significantly, as is the case between platforms supporting multi-qubit gates and those limited to two-qubit gates. In this case, moreover, the fidelity is substantially more costly to calculate. It is also important to note that gate fidelity is not a static property of a quantum hardware. It can vary over time due to coherence times of qubits that are highly sensitive to environmental. Additionally, the fidelity depends on implementation details such as pulse shaping and calibration routines, which can be optimized without manufacture a new quantum hardware. When considering the average gate fidelity of one- and two-qubit gates, which is the metric most commonly used in practice, the metric is practical, as gate fidelities can be efficiently estimated for systems involving at most two qubits. It is also repeatable when evaluated under the same or comparable hardware conditions. However, the metric is not reliable, since it is an average quantity: a higher average gate fidelity does not necessarily imply higher overall computational quality, as a device may achieve high fidelity only for a restricted class of circuits. The metric would be consistent if the average were computed over a common set of generic one- and two-qubit gates. In practice, however, experimental evaluations typically average over native gates, which differ across platforms and architectures. As a result, average gate fidelities cannot be directly compared across different quantum computers. Finally, we do not assess linearity for this metric, as there is no clear notion of how to linearly quantify the overall “quality” of a quantum computer, in contrast to metrics such as execution speed.
[…]
Decoherence times: The decoherence times refer to the period during which the qubit can maintain its quantum coherence. The interaction of qubits with the environment results in the loss of quantum superposition and the decay from the excited state to the ground state, characterized by the times $T_{2}$ and $T_{1}$ respectively. Longer decoherence times are preferable as they might permit deeper circuits to be run, reducing the probability of errors and improving computation accuracy. Additionally, techniques such as dynamical decoupling [134] can be used to improve coherence times by mitigating specific sources of noise applying control pulses to qubits. This metric is practical because $T_{1}$ and $T_{2}$ can be efficiently estimated using standard procedures with low experimental overhead. It is repeatable since repeating the same protocols under similar conditions yields comparable values within statistical fluctuations. It is reliable for characterizing intrinsic coherence limitations, as shorter $T_{1}$ and $T_{2}$ times generally imply faster loss of quantum information and reduced circuit depth capability. However, it is not a complete predictor of quantum processor performance during the execution of a quantum circuit, because gate errors, crosstalk, control imperfections, and measurement errors may dominate even when coherence times are long. It is not a linear metric, since increasing $T_{1}$ or $T_{2}$ does not translate proportionally into higher quantum circuit accuracy or overall quantum processor performance. The metric is consistent because $T_{1}$ and $T_{2}$ have a clear physical meaning and are defined uniformly across platforms.
@@
Gate speeds: Gate speed is a metric that measures how rapidly a quantum computer can execute a given quantum gate [135]. Considering decoherence times, this metric allows us to determine the number of gates that can be performed before the qubits lose coherence. By increasing gate speeds, a greater number of gates can be executed within the same time interval, leading to more applied gates during the lifetime of the qubits. Here, the argument explained for coherence times can be inverted, since fast manipulable systems could also be susceptible to decoherence. The problems of this metric are similar or worse to gate fidelity metric. It is important to highlight that gate speed is not a fixed metric, as it can be optimized through pulse engineering techniques [136]. Moreover, quantum computing platforms, the strength of the interaction between qubits can be dynamically tuned via external control parameters [137]. Therefore, both pulse engineering techniques and dynamic tuning of qubit interactions provide control over quantum gate speeds. This metric is practical, as it only requires executing the quantum gate and measuring its execution time. It is repeatable when the same circuits are run multiple times under comparable experimental conditions. However, it is not reliable, since gate speed is typically evaluated over the native gate set of a quantum computer. Consequently, a device that performs better under this metric does not necessarily execute arbitrary circuits faster than another, as performance depends on how circuits are decomposed into native gates. Moreover, overall performance also depends on qubit coherence times, and faster gates do not necessarily imply better computational performance. For the same reason, the metric is not consistent across different platforms. Finally, the metric is linear: if one system achieves a value that is a factor $x$ smaller than another, it will execute the same circuits in $x$ times less time.